OpenStudy (anonymous):

Write the slope intercept equation for the line with slope -5/6 and y intercept )0,4).... Please show work and GET THE MEDAL

OpenStudy (aroub):

y=mx+b , m= -5/6 , (0,4) 4=-5/6(0)+b 4=b so the equation is y=-5/6 +4

OpenStudy (anonymous):

\[\color{orange}{Same\:\:\:Way\:\:I\:\:Told\:\:You\:\:Before}\]

OpenStudy (anonymous):

ok I got it thanks